OpenStudy (anonymous):

The temperature t in Kelvin is 273.15 more than the temperature c in degrees Celsius. Which function would you use to find the equivalent temperature in Kelvin for 100°C? Find the equivalent temperature

OpenStudy (anonymous):

a. c(t) = t + 273.15; 373.15 K c. c(t) = t + 273.15; 153.75 K b. t(c) = c + 273.15; 373.15 K d. t(c) = c + 273.15; 153.75 K

OpenStudy (aum):

"temperature t in Kelvin is 273.15 MORE than the temperature c in degrees Celsius" implies t = c + 273.15

OpenStudy (aum):

Put c = 100 and find k.

OpenStudy (anonymous):

373.15